Keck Planet Finder (KPF)

Important Notice: KPF will have a “servicing mission” in early 24A to perform several upgrades. This will involve warm up of the optical bench. The expected timeline is to do this in February 2024. Details are still being worked out.

KPF Subsystem Status

A red status indicator denotes a significant problem, while yellow indicates a functional system with limitations, and green indicates a system operating nominally.

Subsystem

Description

Subsystem

Description

Tip Tilt Corrections

The tip tilt stage’s X axis has extremely limited range. This changes with time and is either completely crippling, or completely unnoticeable. This can negatively impact the tip tilt performance in 1 dimension and thus RV precision. As of Nov 30, 2023, the X stage has effectively no range.

Ca HK Detector

Currently non-functional.

Double Star Observations

The tip tilt algorithms have not been optimized for double stars. Observers are welcome to try to observe double stars, but should assume things will not work. Remember that this is not the same as the MAGIQ guiding system, so your intuition there does not transfer to KPF.

Etalon

The etalon is providing substantially less than the expected flux. Simultaneous calibrations are not recommended, but are possible. Slew cals should be taken every hour. We have reasonable flux, but still much less than planned, the problem appears to be a damaged fiber in the etalon. Planning for a long term solution is ongoing.

LFC

Is operating normally. Standard exposure parameters are now ExpTime: 60, CalND1: OD 2.0, and CalND2: OD 0.1.

Detector Systems

The red and green detectors suffer from occasional “start state errors” in which the affected detector remains in the start phase and does not produce a useful exposure. The observing scripts now detect this occurrence, abort the current exposure (with read out) and start a fresh exposure on both cameras. No action is necessary on the part of the observer. This costs about a minute of time for this to happen, but the resulting data should be normal (unless another error occurs). The occurrence rate for these problems is 0.34% on the green detector and 0.22% on the red, so around one in every 180 exposures is affected by one of the two detectors experiencing this error.

Blog Posts

The blog posts for this page will be used to document events which may affect radial velocity measurements of the instrument. The most obvious example are events which induce a significant temperature change in either the detectors or on the spectrograph bench.

 

A fill resulted in a +/- 70 mK temperature transient at the detector. Screenshot 2023-10-30 at 8.52.22 AM.png
A fill problem resulted in a short +50/-60 mK temperature transient. Screenshot 2023-08-15 at 11.16.08 AM.png
Starting on the night of July 27 (HST) the tip tilt stage has had problems with the X axis. This has resulted in poor tip tilt performance. As of Aug 14, troubleshooting is ongoing. Some range was recovered by chance while troubleshooting on 2023-08-24, so the night of 2023-08-25 looked like it had reasonable tip tilt performance. The stage health seems variable, so we are still working on troubleshooting.…
Trouble resetting the detector and reloading the ACF file led to a period of uncontrolled temperatures on Green. See night log K1-33262 for more details. Screenshot 2023-06-28 at 12.49.43 PM.png
The exposure meter DRP has been applying incorrect wavelength solutions to the exposure meter files. This was fixed on or about 2023 June 5 and may have been in place as far back as 2023 Feb 14. The raw 2D files are not saved after the proceeded 1D files are ingested in to KOA, so we are unable to correct old data directly. An example exposure meter spectrum (pulled from Jump) is show below. This spectrum has the bad wavelength solution and the offsets are substantial in the red.…
While work was being done on the green, apparently the red side was disrupted. The archon power appeared off at the power strip (0W being drawn) and no temperatures were being reported. Investigation revealed that the archon switch was on, power cord was in, but chassis was cold (consistent with power strip reading 0W).  The archon box was then power cycled (using switch on archon) and that seems to have brought it up.  Cables run over the green, so they got wiggled during green work,…
The Green detector ceased communicating at about 5am HST. Similar to other events, the CCDPOWER keyword was non responsive, but restarting the service did not help. We power cycled the archon itself from the power strip. Will was able to view camerad logs: poweron ERROR Archon controller returned error processing command: POWERON poweroff ERROR Archon controller returned error processing command: POWEROFF Furthermore, the camerad log (/kroot/var/log/ccd_green/camerad_2023)0519.…
Similar to other recent events http://vm-dashboards.keck.hawaii.edu/d/VoSaQfx4k/kpf-green-thermal-control?orgId=1&from=1683810000000&to=1683824400000 http://vm-dashboards.keck.hawaii.edu/d/VoSaQfx4k/kpf-green-thermal-control?orgId=1&from=1683810000000&to=1683824400000
During troubleshooting of green, we twice set kpfred.CF_BASE_ENA to on as instructed by EMIR notices, but off is apparently the nominal state. This caused temperature excursion, but small as it was caught early. Grafana plot: http://vm-dashboards.keck.hawaii.edu/d/VoSaQfx4l/kpf-red-thermal-control?orgId=1&from=1683763200000&to=1683784800000 http://vm-dashboards.keck.hawaii.edu/d/VoSaQfx4l/kpf-red-thermal-control?orgId=1&from=1683763200000&to=1683784800000 Screenshot 2023-05-10 at 8.18.38 PM.png
Similar to other recent events where CCDPOWER was off. Green side Grafana plot: http://vm-dashboards.keck.hawaii.edu/d/VoSaQfx4k/kpf-green-thermal-control?orgId=1&from=1683763200000&to=1683784800000 http://vm-dashboards.keck.hawaii.edu/d/VoSaQfx4k/kpf-green-thermal-control?orgId=1&from=1683763200000&to=1683784800000 Screenshot 2023-05-10 at 8.14.38 PM.png
Similar to 2023-04-29 at 13:55 HST: KPF Green Detector deltaT=-4.9 we had a temperature excursion when the CCDPOWER keyword got in to a bad state, then temperature set points came back with incorrect values after kpfgreen was restarted. Keck Grafana link: http://vm-dashboards.keck.hawaii.edu/d/VoSaQfx4k/kpf-green-thermal-control?orgId=1&from=1683716400000&to=1683752400000 http://vm-dashboards.keck.hawaii.edu/d/VoSaQfx4k/kpf-green-thermal-control?…
Similar to 2023-04-29 at 13:55 HST: KPF Green Detector deltaT=-4.9 we had a temperature excursion when the CCDPOWER keyword got in to a bad state, then temperature set points came back with incorrect values after kpfgreen was restarted. This time, the difference was noticed (and alarmed on by kpfmon right away, so the excursion was corrected promptly. Keck Grafana instance link to relevant timestamp: http://vm-dashboards.keck.hawaii.edu/d/VoSaQfx4k/kpf-green-thermal-control?…
The green detector went in to an odd state. The first symptom was the inability to take an exposure. A look at keywords revealed that CCDPOWER was not on: 2023-04-19T09:13:41.0630 CCDPOWER = On 2023-04-29T13:55:12.6123 CCDPOWER = Unknown 2023-04-29T13:55:22.6552 CCDPOWER = Not Configured This timestamp (2023-04-29T13:55:12.6123) corresponds to when the detector temperature stopped reporting. Here’s a snapshot of relevant keywords: 2023-04-21T14:12:38.6366 TEMPSET = -140.…
After the Instec pumps had increasing trouble keeping the system cold (the instec cooling power would saturate frequently), impacts began to be seen on the Green detector temperature. The first temperature change on the detector began on 2023-04-18 at 08:27 HST, the system was back under control at about 12:15 HST the same day, but reached a +1.2 degree delta_T. A second temperature excursion began on 2023-04-19 at 11:20 HST and ended at 17:00 HST the same day with a delta_T of +5.4 degrees.…